OpenStudy (anonymous):

Alan is twice as old as Sue and half as old as Joseph. If the average ( arithmetic mean) of all three ages is 14, how old is Sue?

OpenStudy (luke):

a=2s, j=2a, (j+s+a)/3=14. this simplifies to 4s+s+2s=14*3 by using the first two equations to move everything into terms of S and by moving the three over in the euation. this gives s as 6. using the first two equations that gives a as 12 and j as 24.

OpenStudy (anonymous):

Thanx